OpenStudy (anonymous):

Question: How do you turn 1/2(72+x^2)^(-1/2) into x/sqrt(72+x^2)? I get that you flip it, but what happens to the 2nd 1/2 and where does the x in the numerator come from?

OpenStudy (anonymous):

are you taking a derivative? because other wise it is not the same

OpenStudy (anonymous):

ok let start from the beginning

OpenStudy (anonymous):

f(x)=sqrt(72+x^2) = (72+x^2)^(1/2) f'(x)=1/2 (72+x^2)^(-1/2) * (2x) =x / sqrt(72+x^2) f'(7)=7/sqrt(72+49) =7 / sqrt(121) =7 / 11

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

f'(x)=1/2 (72+x^2)^(-1/2) * (2x) =x / sqrt(72+x^2) this is the step im asking about

sam (.sam.):

you were trying to find the derivative of f(x)=sqrt(72+x^2)?

OpenStudy (anonymous):

yes. i just wanna know about that step. everything else makes sense to me and is correct.

sam (.sam.):

\[y=\sqrt{72+x ^{2}}\] \[y=(72+x^2)^{1/2}\] Chain rule, \[y'=(1/2)(72+x^2)^{-1/2}(2x)\] \[y'=\frac{1}{2}\frac{2x}{\sqrt{72+x ^{2}}}\] \[y'=\frac{x}{\sqrt{72+x ^{2}}}\]

sam (.sam.):

you dont understand the chain rule?

OpenStudy (anonymous):

i just dont like the flipping part

OpenStudy (anonymous):

confuses me sometimes

sam (.sam.):